Diagnal brace. The width of a retangular gate is 2 meters (m) larger than its heiht. The diagonal brace measures (Square root) 6 m. Find the width and height.
Answer by edjones(8007)

We have a right triangle of height h, width h+2, and hypotenuse sqrt(6).
a^2+b^2=c^2
h^2+(h+2)^2=sqrt(6)^2
h^2+h^2+4h+4=6
2h^2+4h-2=0
2(h^2+2h-1)=0
h^2+2h =1
h^2+2h+1=1+1 completing the square.
(h+1)^2=2
h+1=+-sqrt(2) Square root of each side.
h= -1+sqrt(2) m
w= 1+sqrt(2) m
.
Ed